OpenStudy (anonymous):

determine that each equation is an identity. (cscX-cotX)^2=1-cosX/1+cosX

OpenStudy (raden):

do for left side, gives : (cscx-cotx)^2 = (1/sinx - cosx/sinx)^2 = ((1-cosx)/sinx)^2 for right side, gives : (1-cosx)/(1+cosx) * (1-cosx)/(1-cosx) = (1-cosx)^2 * (1+cosx)(1-cosx) = (1-cosx)^2 * (1-cos^2 x) = (1-cosx)^2 * sin^2 x = ((1-cosx)/sinx)^2

OpenStudy (raden):

does this make sense ?

OpenStudy (anonymous):

im still figuring it out... but i'm "starting" to understand...

OpenStudy (raden):

oke

OpenStudy (anonymous):

on the right side, why did you do this: (1-cosx)/(1+cosx) * (1-cosx)/(1-cosx)?

OpenStudy (anonymous):

oh never mind i understand :)